# ARCO:Adaptive Multi-Agent Reinforcement Learning-Based Hardware/Software Co-Optimization Compiler for Improved Performance in DNN Accelerator Design

### v2

@@ Abstract.
This paper presents ARCO, an adaptive Multi-Agent Reinforcement Learning (MARL)-based co-optimizing compilation framework designed to enhance the efficiency of mapping machine learning (ML) models - such as Deep Neural Networks (DNNs) - onto diverse hardware platforms. The framework incorporates three specialized actor-critic agents within MARL, each dedicated to a distinct aspect of compilation/optimization at an abstract level: one agent focuses on hardware, while two agents focus on software optimizations. This integration results in a collaborative hardware/software co-optimization strategy that improves the precision and speed of DNN deployments. Concentrating on high-confidence configurations simplifies the search space and delivers superior performance compared to current optimization methods. The ARCO framework surpasses existing leading frameworks, achieving a throughput increase of up to 37.95% while reducing the optimization time by up to 42.2% across various DNNs.
copyright: none
@@ 1. Introduction
The computational demands and complexity of deep neural networks (DNNs) have surged rapidly with their expanding applications across diverse industry sectors, including autonomous driving, medical imaging, and natural language processing (Hossain et al., 2023). These applications require high accuracy and real-time responsiveness, placing significant pressure on the computational infrastructure. Historically, performance optimization in environments like TensorFlow (Abadi et al., 2016) and PyTorch (Paszke et al., 2019) has relied on hand-optimized kernels such as NVIDIA’s cuDNN and Intel’s MKL. While these kernels are designed to optimize efficiency for particular hardware setups, they face challenges in adapting to the changing demands of modern neural architectures and their evolving computational requirements. The extensive number of parameters that need tuning and a huge hardware configuration search space pose significant hurdles (Ahn et al., 2020).
@@
To address this growing complexity, there has been a shift towards automated compilation frameworks like TVM (Chen et al., 2018), TensorComprehensions (Vasilache et al., 2018), and AutoTVM (Chen et al., 2018). These frameworks move away from static, hand-tuned optimization strategies to dynamic, algorithm-driven approaches capable of adapting to various architectures and operational requirements. For example, AutoTVM utilizes a sophisticated approach involving boosted trees (Chen and Guestrin, 2016) to navigate the extensive configuration space of neural network code efficiently. This method predicts the performance of potential configurations and minimizes the need for exhaustive empirical testing, thereby expediting the optimization process. However, despite these technological advances, optimizing complex models requires substantial resources and compilation time, underscoring the ongoing challenges and gaps in current optimization methodologies. In addition to evolutionary strategies like genetic algorithms (Mirjalili and Mirjalili, 2019), machine-learning techniques have been instrumental in refining the search for optimal configurations across this complex landscape. When considering various machine learning algorithms, reinforcement learning (RL) stands out as a promising option for exploring design search spaces and optimizing tasks (Zhang et al., 2022; Wang et al., 2022; Bakshi and Johnsson, 2023; Ahn et al., 2020). RL operates by testing various strategies and utilizing feedback on their actions to maximize cumulative rewards. This learning process occurs through agents interacting with their environment, which can range from single to multi-agent setups. Multi-agent RL offers the advantage of enabling transfer learning and providing valuable insights into interconnected decision-making systems (Du and Ding, 2021).
To address these inefficiencies and push the boundaries of current compilation technologies, we propose an adaptive multi-agent reinforcement learning-based hardware/software co-optimization compiler called ARCO. At its core, this optimizing compiler relies on a multi-agent reinforcement learning (MARL) algorithm to optimize the mapping of the given model and the accelerator structure concurrently. ARCO utilizes three specialized agents—each designed to optimize different facets of the system architecture: two focus on software and DNN configurations, while the third optimizes the accelerator architecture. This actor-critic multi-agent approach facilitates a more integrated and holistic optimization process and allows for more granular control and customization of the optimization strategies employed.
The proposed framework efficiently narrows the vast search space by employing a soft threshold for selecting high-confidence configurations, focusing only on the most promising candidates. This methodology not only reduces computational overhead but also minimizes exploration time. Our method leverages the capabilities of VTA++ (Banerjee et al., 2021), a highly configurable deep learning accelerator, to investigate diverse architectural optimizations. ARCO utilizes Centralized Training with a Decentralized Execution (CTDE) strategy, wherein training occurs in a centralized manner. In contrast, execution is decentralized, enabling adaptive adjustments to the changing requirements of DNN workloads. This strategic approach efficiently narrows the extensive search space by utilizing a soft threshold to select high-confidence configurations, prioritizing the most promising candidates. The efficacy of the ARCO is assessed under a set of DNN models with a variety of architectures and compared with leading state-of-the-art approaches, and it has demonstrated substantial improvements in throughput.
The rest of the paper is organized as follows. Section 2 provides detailed background on the VTA++, CTDE in MARL, workflow for DNN compilers, and related works that have influenced our approach. ARCO and its detailed architecture and components are discussed in Section 3. In Section 4, the efficacy of the ARCO is assessed and compared to that of the established frameworks such as AutoTVM and CHAMELEON (Ahn et al., 2020). Finally, the paper concludes in Section 5.
@@ 2. Background
Figure 1. High-Level architecture of VTA++.
This section briefly introduces the target hardware we used (i.e., VTA++), the Centralized Training and Decentralized Execution in Multi-Agent Reinforcement Learning, the workflow for DNN compilers, and related prior work.
2.1. VTA++
The Versatile Tensor Accelerator (VTA) (Moreau et al., 2019) has emerged as a highly effective deep learning hardware acceleration framework integrated within the TVM stack. To achieve a good hardware/software co-design, we need highly configurable target hardware, which we could optimize its architecture to achieve simultaneous hardware and software optimization. Therefore, we chose VTA++ (Banerjee et al., 2021), which has the same architecture as VTA but is more customizable. The high-level architecture of VTA++ is shown in Fig  1.
The General Matrix Multiply (GEMM) core is the cornerstone of VTA’s capabilities, carrying out dense matrix multiplications. This key operation underpins deep learning functions like 2D convolutions using convolutional layers and classification based on the extracted features using fully connected layers. This core is specifically engineered to handle complex arithmetic calculations by multiplying input and weight tensors and then adding the output to a register file tensor.
The GEMM core of VTA is defined by key hardware parameters: BATCH, BLOCK_IN, and BLOCK_OUT. BATCH size refers to the number of data samples processed in parallel, BLOCK_IN represents the inner dimension of the input tensor, and BLOCK_OUT pertains to the output dimension of the weight tensor and the accumulated result.
In this paper, we dedicate one of our actor-critic MARL agents to optimize the VTA GEMM core for enhanced hardware performance. By selectively tuning BATCH, BLOCK_IN, and BLOCK_OUT—referred to herein as ”hardware knobs”—we adapt the GEMM core to be optimized in our method. Altering BATCH size impacts the parallelism of data processing while modifying BLOCK_IN and BLOCK_OUT influences the granularity of the computation and the utilization of the on-chip memory resources. The goal is to strike an optimal balance between computational throughput and resource allocation, thereby maximizing the GEMM core’s efficiency. This RL agent learns to predict the optimal hardware configuration by interacting with the VTA environment and other software agents, thereby learning from the empirical performance outcomes of previous configurations.
@@ 2.2. CTDE in MARL
Multi-agent reinforcement learning (MARL) extends traditional reinforcement learning, introducing multiple agents within a shared environment. The complexity of agent interactions calls for sophisticated methodologies that coordinate individual agent strategies with collective objectives. The Centralized Training with Decentralized Execution (CTDE) framework has emerged as a significant paradigm in MARL, balancing the collective intelligence during training with the autonomy of individual agents during operation (Lyu et al., 2021).
During the centralized training phase within the CTDE paradigm, agents have access to a global state and learn policies that consider all agents’ collective state and actions. This centralized approach allows agents to develop cohesive strategies, effectively addressing cooperative, collaborative, competitive, and mixed tasks. The essence of CTDE is encapsulated in the adaptation of Proximal Policy Optimization (PPO) (Schulman et al., 2017) to MARL, leading to the development of Multi-Agent PPO (MAPPO) (Yu et al., 2022).
In the context of MAPPO, the learning process involves three pivotal components: critic learning, General Advantage Estimation (GAE), and policy learning. These components are encapsulated in the following mathematical expressions:
  • Critic learning ensures that each iteration yields an enhanced centralized value function. It is defined as:
    (1) $\phi_{k+1}=\arg\min_{\phi}\frac{1}{|D_{k}|T}\sum_{\tau\in D_{k}}\sum_{t=0}^{T}\left(V_{\phi}(\mathbf{o}{t},\mathbf{s}{t},\mathbf{u}{t})-\hat{R}{t}\right)^{2}$
  • General Advantage Estimation (GAE) gauges the quality of the current action relative to the baseline critic value, described by:
@@
  • Policy learning involves computing the policy gradient using the estimated advantage to update the policy function:
@@
    $\text{clip}\left(\frac{\pi_{\theta}(\mathbf{u}|\mathbf{o})}{\pi_{\theta_{k}}},1-\epsilon,1+\epsilon\right)A^{\pi_{\theta_{k}}}(\mathbf{o},\mathbf{s},\mathbf{u}^{\prime})\bigg{)}$
where $D$ represents the trajectories that can be shared across agents, $\hat{R}$ is the estimated reward, $\tau$ denotes a trajectory, and $A$ is the advantage. The discount factor $\gamma$ and the weight of the GAE $\lambda$ govern the temporal scope of the advantage estimation. In these expressions, $\mathbf{u}$ denotes the current agent action, while $\mathbf{u}^{\prime}$ encapsulates the action set of all agents except the acting agent. The global state is represented by $\mathbf{s}$, and $\mathbf{o}$ signifies the local observation of an agent. The hyperparameter $\epsilon$ controls the policy update’s deviation from the previous policy. $V_{\phi}$ is the shared value function, and $\pi_{\theta}$ denotes the shared policy function, highlighting the collaborative essence of CTDE in MARL.
Empirical studies reveal that MAPPO can rival and, at times, outperform competitive off-policy algorithms, which have traditionally been preferred for their sample efficiency (Yu et al., 2022). These findings led us to use MAPPO as our optimization algorithm used in MARL.
@@ 2.3. Workflow for DNN compilers
Mapping DNN models into efficient machine code is crucial for their performance across various hardware platforms. This mapping is managed by a compiler workflow that optimizes the DNN model in several stages (Allen and Kennedy, 2002). Initially, the compiler’s frontend conducts general optimizations. These broad improvements do not yet account for the specific hardware where the model will run. Subsequently, backend optimizations consider the target hardware’s characteristics, shaping the code in a hardware-aware manner. Recent advancements like the AutoTVM system have introduced an innovative step that fine-tunes the code based on actual hardware performance feedback. This method explores various settings—referred to as ”knobs”—to determine the most effective configuration for the code to run efficiently. The search for this optimal configuration ($\Theta^{*}$) is captured by:
@@
where the term $f$ represents the performance or ’fitness’ of the code, and $\Theta$ stands for a specific set of knob settings. The compiler’s goal is to discover the knob settings that maximize the code’s performance on the target hardware within the allowed configuration space $D$.
@@ 2.4. Related Work
In the domain of DNN compilers, adaptive and efficient auto-tuning mechanisms have proven critical for optimizing performance on varying hardware architectures (Rieber et al., 2022; Sadeghi et al., 2024; Ryu et al., 2022; Lu et al., 2022; Fayyazi et al., 2024; Dhakal et al., 2022). An interesting development in this field is AutoTVM (Chen et al., 2018), which utilizes a machine learning-based cost model, specifically XGBoost (Chen and Guestrin, 2016), for its auto-tuning process. This method, while effective, involves a large search space and substantial measurement overhead, making optimization resource-intensive.
@@
MetaTune (Ryu and Sung, 2021) was introduced to address the limitations of black-box auto-tuning by leveraging meta-learning to allow its cost model to ”learn to learn” performance correlations. This enables MetaTune (Ryu and Sung, 2021) to adapt to new optimization spaces quickly and efficiently, demonstrating improvements in inference times compared to prior methods. Glimpse (Ahn et al., 2022) builds upon these approaches by integrating mathematical embeddings of hardware specifications, known as Blueprints, in a Bayesian optimization framework. These Blueprints guide the search algorithm towards subspaces with higher performance potential.
@@
Another approach is CHAMELEON(Ahn et al., 2020), which employs reinforcement learning for its Adaptive Exploration and Sampling algorithms, minimizing invalid configurations and costly hardware measurements. CHAMELEON(Ahn et al., 2020) significantly reduces optimization time compared to AutoTVM while improving inference times. NaaS (Neural Architecture and Accelerator Search) (Zhou et al., 2022) expands upon these frameworks by jointly optimizing neural network architectures and hardware accelerators. It parameterizes both architectures in a unified search space using PyGlove and demonstrates efficiency gains across tasks like image classification and segmentation.
PRIME (Data-Driven Offline Optimization for Architecting Hardware Accelerators) (Kumar et al., 2021) introduces a novel data-driven method for designing hardware accelerators using logged simulation data to construct a robust, conservative surrogate model. By eliminating simulation-driven optimization’s repeated overhead, PRIME significantly reduces simulation time while architecting specialized accelerators for single- and multi-application tasks.
@@
Comparing these methods highlights AutoTVM’s groundwork in machine learning-based auto-tuning. MetaTune and Glimpse advance this by incorporating meta-learning and hardware-aware strategies, respectively. CHAMELEON extends the reinforcement learning approach with improved sampling methods. NaaS and PRIME offer new dimensions by integrating accelerator co-design and offline data-driven optimization. However, despite the advances brought by NaaS and PRIME in hardware consideration, their lack of reinforcement learning makes them slower in compilation times, as they can’t rapidly navigate the expansive hardware and software search space. Reinforcement learning algorithms, as used in CHAMELEON, enable compilers to navigate possible configurations efficiently, identifying high-performance options with fewer tests. This results in a more agile optimization process, allowing for quicker deployment of DNN models and facilitating rapid progress in deep learning research and applications(Zhang et al., 2022; Wang et al., 2022; Bakshi and Johnsson, 2023; Ahn et al., 2020).
@@
In comparing our approach to existing methods like CHAMELEON, Glimpse, and MetaTune—which focus on reducing compilation times through different optimization techniques—and co-design methods like NaaS and PRIME—which strive to enhance system throughput through various optimization strategies—our analysis underscores the distinct advantages of MARL in efficiently navigating the breadth and depth of the design space in modern DNN applications. While effective in their own ways, these methods often fail to fully capture the interdependencies between hardware and software optimizations that our framework does, which can result in suboptimal performance and inefficient throughput in real-world applications. Additionally, co-design optimizers like NaaS and PRIME, though aiming to streamline throughput, lack the capacity to effectively reduce optimization time within the vast hardware-software search space and require significant resources for compiling designs.
3. ARCO
As explained before, current systems fall short of effectively optimizing both the hardware and software components of DNNs simultaneously, as they tend to prioritize software parameters exclusively. Also, existing co-design approaches are too time-consuming to compile. To break free from prolonged optimization processes and fragmented tuning methods, it is important to take on two key challenges: 1) Improving the search algorithm to encompass hardware architecture and parameters as modifiable factors and 2) Refining the sampling process to capture the solution distribution better and minimize encountering non-feasible configurations.
@@ 3. ARCO
We propose two key advancements in the co-optimizing compiler for DNNs to address these obstacles. Firstly, we incorporate a multi-agent reinforcement learning strategy into the search algorithm, allowing it to adapt to new design environments while simultaneously refining hardware and software parameters. This approach harnesses collective intelligence to achieve a more holistic and efficient optimization approach. Secondly, we introduce a novel Confidence Sampling (CS) method to replace the existing uniform (Chen et al., 2018) and adaptive sampling methods (Ahn et al., 2020), which often are random and rely on trial and error. The proposed CS method identifies configurations with a higher probability of success, guided by learned patterns from previous successful optimizations.
3.1. Overview of ARCO
Figure 2. Overall search flow of ARCO.

### v3

# Dynamic Co-Optimization Compiler: Leveraging Multi-Agent Reinforcement Learning for Enhanced DNN Accelerator Performance

@@ Abstract.
This paper introduces a novel Dynamic Co-Optimization Compiler (DCOC), which employs an adaptive Multi-Agent Reinforcement Learning (MARL) framework to enhance the efficiency of mapping machine learning (ML) models, particularly Deep Neural Networks (DNNs), onto diverse hardware platforms. DCOC incorporates three specialized actor-critic agents within MARL, each dedicated to different optimization facets: one for hardware and two for software. This cooperative strategy results in an integrated hardware/software co-optimization approach, improving the precision and speed of DNN deployments. By focusing on high-confidence configurations, DCOC effectively reduces the search space, achieving remarkable performance over existing methods. Our results demonstrate that DCOC enhances throughput by up to 37.95% while reducing optimization time by up to 42.2% across various DNN models, outperforming current state-of-the-art frameworks.
Multi-agaent reinforcement learning, Co-optimization, Hardware accelerators, Sampling, Throughput
journalyear: 2025copyright: rightsretainedconference: 30th Asia and South Pacific Design Automation Conference; January 20–23, 2025; Tokyo, Japanbooktitle: 30th Asia and South Pacific Design Automation Conference (ASPDAC ’25), January 20–23, 2025, Tokyo, Japandoi: 10.1145/3658617.3697547isbn: 979-8-4007-0635-6/25/01ccs: Computing methodologies Multi-agent reinforcement learningccs: Hardware Hardware accelerators
@@ 1. Introduction
The computational demands and complexity of deep neural networks (DNNs) have surged with their expanding applications across various industries such as autonomous driving, medical imaging, and natural language processing (Hossain et al., 2023; Ashrafi et al., 2024; Mokhtari Dowlatabad et al., 2023; Farsi et al., 2024; Razmara et al., 2024). Traditionally, performance optimization in frameworks like TensorFlow (Abadi et al., 2016) and PyTorch (Paszke et al., 2019) has relied on hand-optimized kernels like NVIDIA’s cuDNN and Intel’s MKL. However, these kernels struggle to adapt to the evolving demands of modern neural architectures (Ahn et al., 2020).
@@
To address this, automated compilation frameworks like TVM (Chen et al., 2018), TensorComprehensions (Vasilache et al., 2018), and AutoTVM (Chen et al., 2018) have emerged, shifting from static, hand-tuned optimization to dynamic, algorithm-driven approaches. For instance, AutoTVM uses boosted trees (Chen and Guestrin, 2016) to efficiently explore the configuration space of neural network code. Despite these advances, optimizing complex models still requires substantial resources and compilation time, highlighting ongoing challenges in current optimization methodologies. Reinforcement learning (RL) has shown promise in exploring design search spaces and optimizing tasks (Zhang et al., 2022; Wang et al., 2022; Bakshi and Johnsson, 2023; Ahn et al., 2020), with multi-agent RL offering advantages like transfer learning and insights into interconnected decision-making systems (Du and Ding, 2021).
In this work, we propose an adaptive multi-agent reinforcement learning-based hardware/software co-optimization compiler called DCOC. Our compiler employs a multi-agent reinforcement learning (MARL) algorithm to concurrently optimize the mapping of the DNN model and the accelerator architecture. The use of MARL effectively manages the expansive search space of hardware and software configurations, leveraging the collaborative strengths of multiple agents (Yu et al., 2022), which is crucial due to the significant impact of hardware-software interplay on overall performance.
DCOC features three specialized agents: two focusing on software and DNN configurations, and one optimizing the accelerator architecture. This actor-critic MARL approach enables an integrated and holistic optimization process with granular control over optimization strategies. The framework efficiently narrows the vast search space by employing a soft threshold to select high-confidence configurations, concentrating on the most promising candidates, thus reducing computational overhead and exploration time.
Utilizing a Centralized Training with Decentralized Execution (CTDE) strategy, DCOC allows adaptive adjustments to the changing requirements of DNN workloads. We assess the efficacy of DCOC on a set of DNN models with various architectures, comparing it with leading state-of-the-art approaches. The results demonstrate substantial improvements in throughput, showcasing the potential of MARL-based co-optimization in advancing compilation technologies.
@@ 2.1. CTDE in MARL
Multi-agent reinforcement learning (MARL) extends traditional reinforcement learning by introducing multiple agents within a shared environment, necessitating methods to coordinate strategies towards collective objectives. The Centralized Training with Decentralized Execution (CTDE) framework balances collective intelligence during training with agent autonomy during execution (Lyu et al., 2021).
In CTDE, agents have access to global state information during centralized training, allowing them to learn cohesive policies that consider all agents’ states and actions. This approach effectively addresses cooperative and competitive tasks. Proximal Policy Optimization (PPO) (Schulman et al., 2017) has been adapted to this framework, resulting in Multi-Agent PPO (MAPPO) (Yu et al., 2022).
The MAPPO learning process involves three key components:
  • Critic learning: Enhances the centralized value function:
    (1) $\phi_{k+1}=\arg\min_{\phi}\frac{1}{|D_{k}|T}\sum_{\tau\in D_{k}}\sum_{t=0}^{T}\left(V_{\phi}(\mathbf{o}_{t},\mathbf{s}_{t},\mathbf{u}_{t})-\hat{R}_{t}\right)^{2}$
  • Generalized Advantage Estimation (GAE): Gauges the quality of the current action relative to the baseline:
@@
  • Policy learning: Updates the policy using the estimated advantage:
@@
    $\text{clip}\left(\frac{\pi_{\theta}(\mathbf{u}|\mathbf{o})}{\pi_{\theta_{k}}(\mathbf{u}|\mathbf{o})},1-\epsilon,1+\epsilon\right)A^{\pi_{\theta_{k}}}(\mathbf{o},\mathbf{s},\mathbf{u}^{\prime})\bigg{)}$
Empirical studies show that MAPPO can rival and sometimes outperform competitive off-policy algorithms (Yu et al., 2022), leading us to adopt MAPPO as our optimization algorithm in MARL.
@@ 2.2. Workflow for DNN Compilers
Efficiently compiling DNN models into machine code is crucial for performance across various hardware platforms. Compiler workflows optimize the DNN model in several stages (Allen and Kennedy, 2002). Initially, the compiler’s frontend performs general optimizations without considering specific hardware. Subsequently, backend optimizations tailor the code based on the target hardware’s characteristics.
Recent advancements like AutoTVM have introduced a fine-tuning step that uses hardware performance feedback to explore various configuration settings (”knobs”) and determine the most effective code configuration. The search for the optimal configuration ($\Theta^{*}$) is expressed as:
@@
where $f$ represents the performance metric, $\Theta$ denotes a specific set of knob settings, and $D$ is the allowed configuration space. The compiler’s goal is to discover the settings that maximize performance on the target hardware.
@@ 2.3. Related Work
In the domain of DNN compilers, adaptive and efficient auto-tuning mechanisms are essential for optimizing performance on varying hardware architectures (Rieber et al., 2022; Sadeghi et al., 2024; Ryu et al., 2022; Lu et al., 2022; Fayyazi et al., 2024; Dhakal et al., 2022; Azizi et al., 2024). AutoTVM (Chen et al., 2018) utilizes a machine learning-based cost model, specifically XGBoost (Chen and Guestrin, 2016), but involves a large search space and substantial measurement overhead.
@@
MetaTune (Ryu and Sung, 2021) addresses these limitations by leveraging meta-learning to quickly adapt to new optimization spaces, demonstrating improved inference times. Glimpse (Ahn et al., 2022) integrates mathematical embeddings of hardware specifications, guiding the search algorithm towards higher performance subspaces using Bayesian optimization.
@@
CHAMELEON (Ahn et al., 2020) employs reinforcement learning to minimize invalid configurations and costly hardware measurements, reducing optimization time and improving inference times compared to AutoTVM. NaaS (Zhou et al., 2022) jointly optimizes neural network architectures and hardware accelerators, parameterizing both in a unified search space using PyGlove. PRIME (Kumar et al., 2021) introduces a data-driven method for designing hardware accelerators, using logged simulation data to construct a robust surrogate model, significantly reducing simulation time.
@@
Comparing these methods, AutoTVM lays the groundwork for machine learning-based auto-tuning. MetaTune and Glimpse advance this with meta-learning and hardware-aware strategies, respectively. CHAMELEON enhances the reinforcement learning approach with improved sampling methods. NaaS and PRIME introduce accelerator co-design and offline data-driven optimization but lack reinforcement learning, making them slower in compilation times.
@@
Our approach differs from existing methods like CHAMELEON, Glimpse, and MetaTune—which reduce compilation times through various optimization techniques—and co-design methods like NaaS and PRIME—which enhance system throughput. Our analysis highlights the advantages of MARL in efficiently navigating the design space in modern DNN applications, capturing hardware and software interdependencies to achieve optimal performance and throughput in real-world applications.
3. Proposed DCOC
Existing systems often fail to optimize both hardware and software components of DNNs simultaneously, focusing primarily on software parameters, while co-design approaches are too time-consuming to compile. To address these limitations, we identify two key challenges:
  1. (1)
    Enhancing the search algorithm to include hardware architecture and parameters as adjustable factors.
  2. (2)
    Refining the sampling process to better capture the solution distribution and minimize non-feasible configurations.
@@ 3. Proposed DCOC
We propose two main advancements in the co-optimizing compiler for DNNs. First, we integrate a multi-agent reinforcement learning (MARL) strategy into the search algorithm, enabling simultaneous optimization of hardware and software parameters. This approach leverages collective intelligence for a more holistic and efficient optimization. Second, we introduce a novel Confidence Sampling (CS) method to replace existing uniform (Chen et al., 2018) and adaptive sampling methods (Ahn et al., 2020), identifying configurations with a higher probability of success based on learned patterns.
Our primary optimization goal is to maximize throughput. The search space comprises configurations defined by various knobs, including tile dimensions for batch size, input channels, output channels, and threading options. By systematically exploring these configurations, our framework can adapt to different hardware platforms and optimize performance effectively.
3.1. Overview of DCOC
Figure 1. Overall search flow of DCOC.